OpenStudy (anonymous):

Find the LCM of these numbers. 10, 12, 15 Type the correct answer.

OpenStudy (marihelenh):

60 is the first multiple they have in common. 10 and 15 have 30 in common, but 12 doesn't, so you have to find the next one.

OpenStudy (welshfella):

another way to do this is to write them as prime factors 10 = 2*5 12 = 2*2*3 15 = 3*5 where there are duplicates you just use one of the numbers so LCM = 2*2*3*5 = 60